# Egocentric Scene-aware Human Trajectory Prediction

### v1

@@ 4 Egocentric Navigation Dataset
Figure 5: The dataset contains a mix of weather, road, lighting, and traffic conditions
@@
Our Egocentric Navigation Dataset is collected within a university campus and its vicinity on human participants (under the approval of IRB-60675). The dataset comprises 34 carefully selected collections, each lasting approximately 7 minutes and spanning over 600 meters, designed to capture a wide range of interactions with the environment. These interactions are critical for testing and enhancing trajectory prediction models, particularly in densely populated or complex areas. The dataset is characterized by its diversity, encompassing various weather conditions (rain, sunny, overcast), surface textures (glass, solid, glossy, reflective, water), and environmental features (stairs, ramps, flat grounds, hills, off-road paths), alongside dynamic obstacles, including humans. Such variety ensures the dataset’s applicability across a wide spectrum of egocentric navigation and prediction tasks.
@@
Recorded at a 20Hz sampling rate, the dataset includes comprehensive state and visual information to capture the nuances in human behavior: 6 degrees of freedom (dof) torso pose in a global frame, leg joint angles (hips and knees of both legs), torso linear and angular velocity, and gait frequency. Visual data comprise aligned color and depth images, semantic segmentation masks, and visual memory frames generated to aid in trajectory prediction. This extensive collection amounts to 198 minutes of data (over 400GB). In practice, we find it is possible to train a very high-quality model even with a smaller dataset size. Therefore we curated a high-quality pilot dataset with roughly 15% of the full data. This allows us to quickly iterate with faster training and an acceptable amount of performance degradation. The performance has been qualitatively compared in Sec. 5.
@@
Acknowledging the limitations of existing datasets in this domain, such as the TISS dataset[11], our Egocentric Navigation Dataset addresses critical gaps by providing dense, high-frequency logs with rich visual and state information. Previous datasets have often suffered from sparse content, low logging frequencies, and lack of access to raw data or pre-trained decoder weights, hindering the development of sophisticated models and methods. To foster innovation and accessibility in egocentric trajectory prediction research, we are committed to open-sourcing our dataset following the de-identification of all faces within the data. We believe this approach will democratize access to high-quality data, enabling researchers to explore new methodologies and applications in the field.
5 Evaluation
5.1 Hardware Setup
We followed the hardware setup illustrated in this previous work [24]. To accommodate data collection in both indoor and outdoor environments, we opted for a setup that does not rely on GPS due to its indoor limitations and avoids IMU localization to minimize drift over extended periods. For precise localization, we employ an Intel Realsense T265 for SLAM-based tracking. We upgraded the stereo camera to an Intel Realsense D455, benefiting from its longer stereo baseline to enhance depth perception and extend the observable range. Considering the variability in lighting conditions, especially outdoors, we chose not to use LiDAR, despite its accuracy in optimal settings. The entire software stack, including real-time panorama generation and data storage, and physically powering the sensor is handled by an Apple Silicon MacBook Pro housed in a backpack. This setup ensures mobility and flexibility during data collection.
5.2 Evaluation Metrics
@@
To comprehensively evaluate the predicted trajectories, we introduce three key metrics: collision-free score, smoothness, and best of N (including best of 1 as a subset for comparison with unimodal predictions). These metrics are designed to reflect the multifaceted nature of trajectory prediction, capturing aspects of physical feasibility, motion quality, and alignment with potential human preferences. An infographic summarizing these metrics is depicted in Fig. 6.
5.2.1 Collision-free Score
@@
This metric assesses the feasibility of the predicted trajectories by ensuring they do not intersect with impermeable objects within the environment. Given the absence of ground truth environmental meshes, we evaluate collisions against a discrete point cloud, defining a collision as the presence of more than 10 scene points within a 0.16m radius of the predicted position. The collision-free score tallies the number of consecutive prediction steps where the trajectory avoids collision with static obstacles such as the ground, stairs, walls, and rough terrains, excluding doors and movable objects. A higher score, up to the total number of prediction steps, indicates better performance.
5.2.2 Smoothness
$Smoothness=1/(\frac{\sum^{n}_{i=1}|V_{i}-\hat{V}_{i}|}{n}+\frac{\sum^{n}_{i=1}|a_{i}-\hat{a}_{i}|}{n})$ (2)
5.2.3 Best of N & Best of 1
Recognizing the inherent multiplicity of valid trajectories for any given scenario, the Best of N metric evaluates the model’s ability to encapsulate the ground truth trajectory within its distribution of predictions. This approach mirrors the minADE-K metric used in autonomous vehicle research, focusing on the minimum average displacement error across a set of K predictions. The Best of 1 metric complements this by assessing the accuracy of a single prediction, facilitating comparison with models that do not generate a multimodal distribution. It also shows how well the estimated distribution aligns with the ground truth under the assumption that the ground truth mode should have the highest likelihood in the given scenario. Together, these metrics ensure a balanced evaluation of the model’s predictive capacity and its versatility in capturing the range of plausible ego-motions.
Figure 7: BEV: Typical input and output are visualized top-down. Five trajectories are predicted in this case, each with a different shade of red. Visual memory is projected to a point cloud and color-coded same as semantic.
5.3 Prediction Evaluation
Our evaluation focuses on assessing the model’s capacity for future trajectory prediction and its ability to generate expected future visual memory encodings, as detailed in Sec. 3. Operating over a 100-step (5-second) prediction horizon—aligned with the typical visibility range (8-meter radius) and walking speeds—our method ensures the generation of predictions that are both relevant and plausible within the observable environment. The inclusion of future visual memory alongside trajectory predictions enhances our ability to visually validate the model’s anticipations of scene evolution, providing a robust cross reference for trajectory prediction and insight into the model’s typical failure modes.
A pertinent example of this evaluation is depicted in Fig. 8, where visual memories associated with the red trajectory prediction reveal a challenge commonly encountered by generative models: the occasional "hallucination" of non-existent features in scenarios of occluded vision or gaps in the observable scene. For instance, as a person navigates a hallway, gaps in conditioning visual memories might occur due to changes in view angle or partial occlusions. At a specific timestep (T3), the model inaccurately anticipates that approaching a gap would reveal a larger, unexplored space (as in T6), leading it to "hallucinate" an extension of the hallway based on patterns observed in the training data. This results in the model predicting a trajectory that erroneously veers into this fictitious extension (T8), as evidenced by the gradual shift of the actual hallway from the center of the predicted path.
Figure 8: Analysis of fail prediction: A failed prediction (bright red) is picked to analyze the failure mode. Predicted visual memory frames are sampled along the trajectory to show the model’s expectation. Visual memory is always centered to forward direction, turning can be observed by following the centerline. BEV plot: Blue trajectory is past, and green is the ground truth. Point clouds follow semantic color codes.
5.4 Ablations
To test the effectiveness of individual modules proposed in our method, we conducted the following ablation studies. The individual ablations and corresponding metrics are summarized in Tab. 1. As mentioned in Sec. 4, we have a full dataset for the best possible performance and a curated pilot dataset for quick iteration. All of the ablation training, unless specifically mentioned, are trained on the pilot dataset. The entries in the table have been sorted by the collision-free score in the second column. The ground truth trajectory did not receive a perfect score (99) in collision avoidance, due to the inaccuracies in the SLAM pose estimation and depth artifacts from the stereo camera. In rare occasions, ground truth trajectory collides with a part of the point cloud that appears a considerable distance away from the ground truth surface of the scene.
Scalability To quantify the difference between training on a large dataset and the pilot dataset, we trained two versions of the baseline one on each dataset. The full dataset model is better on all of the evaluation metrics as expected. However the gap is quite small on the collision metric, we believe the collision avoidance performance is highly dependent on the accuracy of the point cloud, otherwise the model would think that the ground truth demonstration has penetrated some points despite the points being from artifacts and misestimations. On the other hand, the added samples have greatly improved the variance of the generation and hence better Best of n scores in general.
Transformer UNet vs UNet Adding in the multi-head self-attention layers (MHSA) between down and up blocks in UNet significantly blows up the size of the model, fortunately, the inference and training speeds have remained largely unchanged. Without attention layers, the most significant degradation in performance is in collision avoidance. This corroborates our expectation that the MHSA layers help to relate encoded visual memory with trajectory samples.
Hybrid Generation As proposed in Sec. 3, hybrid generation can greatly speed up generation at the expense of a slight decrease in quality (as quantified in ablation), up to 50 trajectory predictions per second on a single Nvidia A5000 GPU. In this experiment, our hybrid generation uses 20 steps of DDIM, followed by 10 steps of DDPM to refine the generation. Compared to a full 1000-step DDPM baseline, this method is 33 times faster, but the quantitative metrics only decreased slightly. To show the effectiveness, we also include the same 30-step generation result except from DDIM. Both collision and smoothness scores were negatively impacted especially the smoothness score. The best-of-n metrics remain similar, showing that while the DDIM is converging in the right direction, it does not converge to a strong solution like DDPM.
Past Condition Length To achieve the best possible performance, we are using a non-Markovian assumption when performing prediction. This requires us to provide a long history of past trajectories. We are curious as to how well the Markovian assumption will hold for this task and are interested in potentially reducing the number of inputs to the model for more efficient use of computation. Therefore we trained a variant that only takes in 3 steps of past trajectory steps, essentially only providing information up to the order of acceleration. The resulting model did not suffer too much from the change, showing that the Markovian assumption largely holds for most of the scenarios.
Semantic Segmentation To test the effectiveness of semantic segmentation. The VAE and diffusion model are trained on the same dataset except without a semantic channel. We see a significant drop in collision avoidance performance. This stems from the same origin as the quality issue we mentioned in Scalability above. Without semantics, the model will not be able to differentiate between penetrating the door and the wall and, thus tends to ignore the geometric constraints provided in the visual memory.
Visual Input Empirically, the removal of the visual input results in the worst-performing model. This emphasizes again the effectiveness of contextual structure provided by the visual memory. Upon inspecting the results, many of the generated samples are simply memorizing what was in the training set and the distribution is incorrect (which also corresponds to a high best-of-1 score as mentioned in Evaluation Metrics of Sec. 5).
@@ 5.4 Ablations
Table 1: Metric comparison table across ablations
Ablation | Collision ↑| Smoothness ↑| Best of 1 ↓| Best of 15↓
@@
w/ Hybrid generation (I20 P10) 88.7 2.17 0.89 0.49
@@
5.5 Comparison with Baselines
The proposed method is compared to existing methods LSTM and CXA Transformer. For consistency, all three methods are based on the same pre-trained VAE. And they are all trained and evaluated on the pilot dataset. Results are shown in Tab. 2. We would like to highlight that some key differences between our model and the baselines are: 1) Our model does not rely on auto-regressive generation and, therefore less affected by compounding error problems. Our model also generally requires less pass through the generation model, which means faster speed. 2) Our model is inherently based on a random process, therefore it learns an implicit distribution of preferences instead of simply an average.
VAE-LSTM One of the most common paradigms to the trajectory prediction task in the existing literature [24, 11] is to use an encoder head to understand the visual input, and then pass the state information along with encoded visual cues into a recurrent model to auto-regressively predict the future states. The quality of the output is usually pretty good. The biggest drawback of this approach is that the prediction is deterministic, as the variance only exists in the encoding process of the visual input, not in the trajectory latent space. And since there is only one deterministic prediction in each scenario, there is zero guarantee of the temporal consistency of the prediction between consecutive scenarios. Despite the drawbacks, the auto-regressive model excels at generating smooth motion as every step is based on the previous predictions, which is also confirmed by the smoothness score. We implemented the LSTM baseline according to the receipt provided in this egocentric trajectory prediction paper[24]. The LSTM model in the baseline performs surprisingly well at avoiding obstacles, given its tiny size of fewer than 1 million parameters.
CXA-Transformer One of the newer approaches is to use the transformer instead of the recurrent modules[11]. The transformer models, at its core, rely on attention mechanisms to process the sequence so it suffers less from forgetting and improves computational efficiency for scaling. One of the most recent works aiming to tackle egocentric trajectory prediction in a similar setup proposes a Cascaded Cross-Attention(CXA) Transformer. It uses a transformer encoder for each modality of input including surrounding people’s pose, semantic segmentation, and past trajectories, and a cross-attention module to fuse in multi-modal information. Similar to other transformer models, it uses a transformer decoder to produce a prediction of the most likely future path auto-regressively. CXA transformer is implemented strictly as outlined and uses hyper-parameters in the paper. In training and evaluation, we noticed a strong tendency to overfit to the training set, despite adding a significant amount of dropouts and many hyper-parameter tuning. The best result we can get is shown in the table.
@@ 5.5 Comparison with Baselines
Table 2: Metric comparison table across baselines
Model | Collision ↑ | Smoothness ↑ | Best of 1 ↓ | Best of 15 ↓
@@
6 Conclusion
This paper presented a pipeline to predict future trajectories using diffusion model conditioned on the past trajectory and scene semantics. We show that it benefits from semantic segmentation, and outperforms existing methods. Limitations of our method include the current assumption that the scene is static, despite it is implicitly handled by semantic segmentation, it can still get quite challenging when there are numerous people around.

### v3

@@ IV Egocentric Navigation Dataset
Figure 6: Dataset: The dataset contains a healthy mix of weather, road, lighting, and traffic conditions
@@
Our Egocentric Navigation Dataset is collected within a university campus and its vicinity on human participants (under the approval of IRB-60675). The dataset comprises 34 carefully selected collections, each lasting approximately 7 minutes and spanning over 600 meters, designed to capture a wide range of interactions with the environment. These interactions are critical for testing and enhancing trajectory prediction models, particularly in densely populated or complex areas. The dataset is characterized by its diversity, encompassing various weather conditions (rain, sunny, overcast), surface textures (glass, solid, glossy, reflective, water), and environmental features (stairs, ramps, flat grounds, hills, off-road paths), alongside dynamic obstacles, including humans. Such variety ensures the dataset’s applicability across a wide spectrum of egocentric navigation and prediction tasks. It is worth noticing that we did not deliberately avoid any pedestrians, reflecting real-world campus walking scenarios.
@@
Recorded at a 20Hz sampling rate, the dataset includes comprehensive state and visual information to capture the nuances in human behavior: 6 degrees of freedom (dof) torso pose in a global frame, leg joint angles (hips and knees of both legs), torso linear and angular velocity, and gait frequency. Visual data comprise aligned color and depth images, semantic segmentation masks, and visual memory frames generated to aid in trajectory prediction. This extensive collection amounts to 198 minutes of data (over 400GB). In practice, we find it is possible to train a very high-quality model even with a smaller dataset size. Therefore we curated a high-quality pilot dataset with roughly 15% of the full data. This allows us to quickly iterate with faster training and an acceptable amount of performance degradation. The performance has been qualitatively compared in Sec V.
Previous datasets have often suffered from sparse content, low logging frequencies, and lack of access to raw data or pre-trained decoder weights, hindering the development of sophisticated models and methods. Acknowledging the limitations of existing datasets in this domain, such as the TISS dataset[31], our Egocentric Navigation Dataset addresses critical gaps by providing dense, high-frequency logs with rich visual and state information. Recently, Meta released the Aria Everyday Activities Dataset (AEA)[40], which also adopts an ego-centric view. Our dataset offers distinct advantages for outdoor navigation scenarios by providing dense depth information from sensors. In contrast, AEA focuses on indoor scenes and daily manipulation activities, offering SLAM key points rather than dense depth maps. While both datasets contribute to egocentric research, ours is particularly well-suited for tasks requiring detailed depth perception and outdoor navigation, allowing researchers to choose the most appropriate resource for their specific research needs.
@@
To foster innovation and accessibility in egocentric trajectory prediction research, we are committed to open-sourcing our dataset following the de-identification of all faces within the data. We believe this approach will democratize access to high-quality data, enabling researchers to explore new methodologies and applications in the field.
@@
IV-A Dataset Detail
All the data are collected at a rate of 20Hz, both state and visual inputs. State information includes 6 dof torso localization in the global frame, leg joint angles (left, right calf and thigh), torso velocity, torso angular velocity, and average gait frequency in 2 seconds moving window. Visual information includes RGB images, an aligned depth frame from D455 at a resolution of 848x480, semantic segmentation masks by DINOv2 + Mask2Former segmentation head, as well as panoramas with all the aforementioned channels projected to 360 view. This brings the total dataset to 198 minutes, and over 400GB.
We optimize storage by retaining raw depth frames only when there’s a significant change in camera position—specifically, an angular movement exceeding 15 degrees or a translation beyond 1 meter. This strategy reduces the point cloud processing rate to approximately 1 keyframe per second at normal walking speeds, thus conserving computational resources for semantic segmentation. Visual memories are generated in a separate process at 20Hz, center-aligned with the current camera pose, to support timely model inference.
We found that the raw data from stereo cameras requires prepossessing to enhance quality, as it tends to connect disjoint edges in the scene. We apply a tuned Canny edge filter to mitigate artifacts along object edges, removing up to 10 pixels around the disjoint edges, this prepares the depth frames for more accurate panorama construction.
We define eight semantic classes for the segmentation: No label, normal ground, stair, door, wall, obstacle, movable, and rough ground. The collision evaluation framework specifically considers ground, stairs, walls, obstacles, and rough terrains to accurately differentiate between navigable and non-navigable spaces.
V Evaluation
V-A Hardware Setup
We utilized a similar hardware setup as this previous work [9] and made some upgrades. To accommodate data collection in both indoor and outdoor environments, we opted for a setup that does not rely on GPS due to its indoor limitations and avoids pure IMU localization to minimize drift over extended periods. For precise localization, we employ an Intel Realsense T265 for SLAM-based tracking. We upgraded the stereo camera to an Intel Realsense D455, benefiting from its longer stereo baseline to enhance depth perception accuracy and extended the observable range. Considering the variability in lighting conditions, especially outdoors, we chose not to use LiDAR, despite its accuracy in optimal lighting conditions. The entire software stack, including real-time panorama generation and data storage, and physically powering the sensor is handled by an Apple Silicon MacBook Pro housed in a backpack. This setup maximizes mobility and flexibility during the data collection.
V-B Evaluation Metrics
@@
To comprehensively evaluate the predicted trajectories, we introduce three key metrics: collision-free score, smoothness, and best of N (including best of 1 as a subset for comparison with unimodal predictions). These metrics are designed to reflect the multifaceted nature of trajectory prediction, capturing aspects of physical feasibility, motion quality, and alignment with potential human preferences. An infographic summarizing these metrics is depicted in Fig 7.
V-B1 Collision-free Score
@@
This metric assesses the feasibility of the predicted trajectories by ensuring they do not intersect with impermeable objects within the environment. Due to the absence of ground truth environmental meshes, we evaluate collisions against a discrete point cloud. To calculate the collision-free metric, we first re-project the visual memory back into a 3D point cloud. At each predicted time step, we retrieve the closest 20 points in the point cloud using a K-D Tree. We define a collision as the presence of more than 10 scene points within a 16cm radius of the predicted position. If a collision is detected, the current and all subsequent time steps are deemed collided. The collision-free score tallies the number of consecutive prediction steps where the trajectory avoids collision with static obstacles such as the ground, stairs, walls, and rough terrains, excluding doors and movable objects. The metric value is the index of the first collided time step, with a maximum score of n-1 for n prediction steps. A higher score indicates better performance. The radius (16cm) and threshold (10 points) parameters are tunable, and these empirical values were found to be most reasonable for our resolution.
@@
TABLE I: Metric comparison table across ablations. Bold = Best, Italic = 2nd
Ablation Collision ↑ Smoothness ↑ Best of 1 ↓ Best of 15↓
@@
w/ Hybrid generation (DDIM20+DDPM10) 88.7 2.17 0.89 0.49
@@
V-B2 Smoothness
$Smoothness=1/(\frac{\sum^{n}_{i=1}|V_{i}-\hat{V}_{i}|}{n}+\frac{\sum^{n}_{i=1}|a_{i}-\hat{a}_{i}|}{n})$
V-B3 Best of N & Best of 1
Recognizing the inherent multiplicity of valid trajectories for any given scenario, the Best of N metric evaluates the model’s ability to encapsulate the ground truth trajectory within its distribution of predictions. This metric is also known as the minADE-K metric widely used in trajectory prediction tasks especially in the autonomous driving field, focusing on the minimum average displacement error across a set of K predictions. The Best of 1 metric, equivalently the ADE metric, complements this by assessing the accuracy of a single prediction, facilitating comparison with models that do not generate a multimodal distribution. It also shows how well the estimated distribution aligns with the ground truth under the assumption that the ground truth mode should have the highest likelihood in the given scenario. Together, these metrics ensure a balanced evaluation of the model’s predictive capacity and its versatility in capturing the range of plausible ego-motions.
Figure 8: BEV: Typical input and output are visualized top-down. Five trajectories are predicted, each with a different shade of red. Visual memory is projected to a point cloud and color-coded same as semantic.
V-C Prediction Evaluation
Our evaluation focuses on assessing the model’s capacity for future trajectory prediction, and its ability to generate expected future visual memory encodings, as detailed in Sec III. Operating over a 100-step (5-second) prediction horizon—aligned with the typical visibility range (8-meter radius) and walking speeds—our method ensures the generation of predictions that are both relevant and plausible within the observable environment. An example of output is visualized in Fig 8. The figure shows everything in a top-down view, specifically the visual memory given is turned into a point cloud with different colors representing different semantic classes. We can see a wall on the top and two pillars on the bottom, along with some people. The 5 predictions from the model are labeled in red, and the actual ground truth is green. Upon visual inspection, the outputs are diverse and all feasible given the scenario.
The inclusion of future visual memory alongside trajectory predictions enhances our ability to visually validate the model’s anticipations of scene evolution, providing a robust cross reference for trajectory prediction and insight into the model’s typical failure modes. A pertinent example of this evaluation is depicted in Fig 9, where visual memories associated with the red trajectory prediction reveal a challenge commonly encountered by generative models: the occasional ”hallucination” of non-existent features in scenarios of occluded vision or gaps in the observable scene. For instance, as a person navigates a hallway, gaps in conditioning visual memories might occur due to changes in view angle or partial occlusions. At a specific timestep (T3), the model inaccurately anticipates that approaching a gap would reveal a larger, unexplored space (as in T6), leading it to ”hallucinate” an extension of the hallway based on patterns observed in the training data. This results in the model predicting a trajectory that erroneously veers into this fictitious extension (T8), as evidenced by the gradual shift of the actual hallway from the center of the predicted path.
Figure 9: Analysis of fail prediction: A failed prediction (bright red) is picked to analyze the failure mode. Predicted visual memory frames are sampled along the trajectory to show the model’s expectation. Visual memory is always center aligned to the forward direction, turning can be observed by following the centerline. BEV plot: The blue trajectory is the past, and the green is the ground truth. Point clouds follow semantic color codes.
V-D Ablations
To test the effectiveness of individual modules proposed in our method, we conducted the following ablation studies. The individual ablations and corresponding metrics are summarized in Table I. As mentioned in Sec IV, we have a full dataset for the best possible performance and a curated pilot dataset for quick iteration. All of the ablation training, unless specifically mentioned, are trained on the pilot dataset. The entries in the table have been sorted by the collision-free score in the second column. Notice that the ground truth trajectory did not receive a perfect score (99) in collision avoidance, this is due to the inaccuracies in the SLAM pose estimation and depth artifacts from the stereo camera. In rare occasions, ground truth trajectory collides with a part of the point cloud that appears a considerable distance away from the ground truth surface of the scene.
Semantic Segmentation In this task, VAE and diffusion model are trained on the same dataset except without the semantic channel. We see a significant drop in collision avoidance (-5.1). Without semantics, the model will not be able to differentiate between penetrating the door vs the wall. This conflicting behavior in training data makes the model tends to ignore the geometric constraints provided in the visual memory.
Visual Input As expected, the complete removal of the visual input results in the worst-performing model. This emphasizes again the effectiveness of contextual structure provided by the visual memory. Upon inspecting the results, many of the generated samples are rote memorization of the training set and the distribution is largely incorrect (which corresponds to the worst best-of-1 score as explained in Evaluation Metrics of Sec V).
Scalability To quantify the difference between training on a large dataset and the pilot dataset, we trained two versions of the baseline one on each dataset. The full dataset model is better on all of the evaluation metrics as expected. However the gap is quite small on the collision metric (+0.6), we believe the collision avoidance performance is more sensitive to the accuracy of the point cloud, than the amount of the training data. If the point cloud quality is low, the model would think that the ground truth demonstration has penetrated some points despite the points are from artifacts and mis-estimations. On the other hand, the added samples have greatly improved the variance of the generation and hence better mode coverage (Best of n -0.06) and best sample quality (Smoothness +2.72)
Transformer UNet vs UNet Adding in the multi-head self-attention layers (MHSA) between down and up blocks in UNet significantly blows up the size of the model, fortunately, the inference and training speeds have remained largely unchanged. With attention layers, the most significant boost in performance is in collision avoidance (+4.0). This corroborates our expectation that the MHSA layers help to relate encoded visual memory with trajectory samples.
@@ V-D Ablations
TABLE II: Metric comparison table across baselines. Bold = Best, Italic = 2nd
Model Collision ↑ Smoothness ↑ Best of 1 ↓ Best of 15 ↓
@@
Hybrid Generation & Computation Efficiency As proposed in Sec III, hybrid generation can greatly speed up generation at the expense of a slight decrease in quality. We were able to obtain 50 trajectory predictions per second on a single Nvidia A5000 GPU. In practice, this is enough to give a sizeable distribution covering the majority of the possible cases in the scene. The GPU utilization rate is usually above 95% during inference. In this experiment, our hybrid generation uses 20 steps of DDIM, followed by 10 steps of DDPM to refine the generation. Compared to a full 1000-step DDPM baseline, this method is 33 times faster, but the quantitative metrics only decreased slightly (Collision -0.5, Best of n +0.02). To show the effectiveness, we also include the same 30-step generation result except for all steps from DDIM. Both collision (-3.9) and smoothness scores (-1.58) were both significantly negatively impacted. This shows that while DDIM is converging in the right direction, it does not converge to as strong a solution as DDPM.
Past Condition Length To achieve the best possible performance, we are using a non-Markovian assumption when performing the prediction, meaning a long history of past trajectories is provided. We are curious as to how well the Markovian assumption holds and are interested in potentially reducing the number of inputs to the model for more efficient use of computation. Therefore we trained a variant that only takes in 3 steps of past trajectory steps, essentially only providing information up to the order of acceleration. The result: Collision only slightly decrease (-0.4), since collision performance is mostly based on visual information; Smoothness is notably worse than before (-0.48), as the model receives less information about the momentum; Mode coverage is impacted slightly (Best of N +0.05). This shows that the Markovian assumption largely holds for most of the scenarios. And we can potentially speed up more using this model in deployment.
V-E Comparison with Baselines
The proposed method is compared to the current state-of-the-art methods: LSTM and CXA Transformer. For consistency, all three methods are based on the same pre-trained VAE. And they are all trained and evaluated on the pilot dataset. Results are compiled in Table II. We would like to highlight some key differences between our model and the baselines: 1) Our model does not rely on auto-regressive generation and, therefore less affected by compounding error problems. Thus also requires less pass through the generation model, meaning faster speed. 2) Our model is inherently based on a random process, therefore it learns an implicit distribution of preferences instead of simply an average.
VAE-LSTM One of the most common paradigms to the trajectory prediction task in the existing literature [9, 17, 21, 25, 12] is to use an encoder head to understand the visual input, and then pass the state information along with encoded visual cues into a recurrent model to auto-regressively predict the future states. The biggest drawback of this approach is that the prediction is deterministic, as the variance only exists in the encoding process of the visual input, not in the trajectory latent space. And since there is only one deterministic prediction in each scenario, there is zero guarantee of the temporal consistency of the prediction between consecutive time steps. Despite the drawbacks, the auto-regressive model excels at generating smooth motion as every step is based on the previous predictions, which is also confirmed by the smoothness score (+7.05). The LSTM baseline performs surprisingly well at avoiding obstacles at a tiny size of fewer than 1 million parameters.
CXA-Transformer The more recent approach is to use the transformer instead of the recurrent modules[31, 18]. The transformer models, at its core, rely on attention mechanisms to process the sequence so it suffers less from forgetting and improves computational efficiency for scaling. One of the most recent works[31] aiming to tackle egocentric trajectory prediction proposes a Cascaded Cross-Attention(CXA) Transformer. It uses a transformer encoder for each modality of input including surrounding people’s pose, semantic segmentation, and past trajectories, and a cross-attention module to fuse in multi-modal information. Similar to other transformer models, it uses a transformer decoder to produce a prediction of the most likely future path auto-regressively. CXA transformer is implemented strictly as outlined in the paper. To optimize its performance on our dataset, we started from the provided hyper parameters and did some tuning. Notably, the collision score of the CXA transformer is significantly lower (-8.9), potentially because it uses stereo camera frames directly with a limited field of view. Unfortunately, the CXA transformer also only returns a single modal prediction. Despite that, the Best of 1 score is still worse than our baseline (+0.38). As an auto-regressive model, the smoothness is still its strength (+1.66).
This showcases the unique advantage of our method using the diffusion model for multi-modal prediction, and Visual Memory to maximize the situational awareness of the model.
VI Conclusion
This paper presented a novel pipeline to predict future trajectories using a diffusion model conditioned on the past trajectory and scene semantics. In the process, we proposed an efficient representation called Visual Memory, a new sampling scheme to speed up sample generation for real-time prediction, and a diverse egocentric human walking dataset to enhance future research in the field. We showed that our method is highly effective through comprehensive ablations and comparisons with current state-of-the-art baselines.